OpenStudy (yanni):

-3(3v-2)+6v=7+6(2v-1) simplify

sam (.sam.):

-3(3v-2)+6v=7+6(2v-1) -9v+6+6v=7+12v-6 -3v+6=12v+1 15v=5 v=1/3

OpenStudy (yanni):

thx